OpenStudy (anonymous):

Prove: cos θ - cos θ•sin^2 θ = cos^3 θ

OpenStudy (amoodarya):

first factor cos \[\cos(\theta)-\cos(\theta)\sin^2(\theta)=\\cos(\theta)(1-\sin^2(\theta))=\\\] then use : \[1-\sin^(\theta)=\cos^(\theta)\]